In an experiment to determine the enthalpy of solution of ethylenediamine, 50.0mL of 2.00M ethylenediamine was added to 50.0mL of water in a coffee cup at 25.00°C. The temperature increased to a final temperature of 27.25°C. Assuming that the heat capacity of the coffee cup is 46.8J/°C, calculate the enthalpy of solution of ethylenediamine. You may assume that all aqueous solutions have a density of 1.00g/mL and a specific heat of 4. 1 84J/g°C. 
Answer #1

moles of ethylene diamine = 50 x 2.0 / 1000 = 0.1

total volume of solution = 50 mL + 50 mL =100 mL

density = 1 g / mL

mass of solution = 100 x 1 = 100 g

dT = 27.25 - 25.00 = 2.25 oC

Q = m Cp dT + Cp dT

Q = 100 x 4.184 x 2.25 + 46.8 x 2.25

Q   = 1046.7 J = 1.05 kJ

enthalpy of solution = - Q / n

                               = -1.05 / 0.1

                              = -10.5 kJ /mol

enthalpy of solution = -10.5 kJ / mol